OpenStudy (yusmin):

y=-x-7 (complete the order pairs) (-5,?) (-2,?) please help

jimthompson5910 (jim_thompson5910):

Ordered pairs are of the form (x,y) x always comes first before y When they say (-5, ?) they mean that x = -5 and y is unknown

jimthompson5910 (jim_thompson5910):

what you do is replace every x with -5 and simplify to find y y = -x-7 y = -1*x-7 y = -1*(x)-7 y = -1*(-5)-7 ... replace every x with -5 y = 5-7 y = -2 So when x = -5, the value of y is y = -2

jimthompson5910 (jim_thompson5910):

so (-5, ?) becomes (-5, -2)

OpenStudy (yusmin):

ohhhhhhhh

jimthompson5910 (jim_thompson5910):

if x = -2, then what must y be?

OpenStudy (yusmin):

some reason I got -7,1

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (yusmin):

because I have no idea how to really do it and I never replaced anything

jimthompson5910 (jim_thompson5910):

I see

jimthompson5910 (jim_thompson5910):

if x = -2, then what must y be? Use the similar steps I posted

OpenStudy (yusmin):

-5? @jim_thompson5910